### v3

# How to Adapt Control Barrier Functions? A Learning-Based Approach with Applications to a VTOL Quadplane

Abstract
In this paper, we present a novel theoretical framework for online adaptation of Control Barrier Function (CBF) parameters, i.e., of the class $\mathcal{K}$ functions included in the CBF condition, under input constraints. We introduce the concept of locally validated CBF parameters, which are adapted online to guarantee finite-horizon safety, based on conditions derived from Nagumo’s theorem and tangent cone analysis. To identify these parameters online, we integrate a learning-based approach with an uncertainty-aware verification process that account for both epistemic and aleatoric uncertainties inherent in neural network predictions. Our method is demonstrated on a VTOL quadplane model during challenging transition and landing maneuvers, showcasing enhanced performance while maintaining safety. [Project Page] [^1] Project page: https://www.taekyung.me/how-to-adapt-cbf[Code (VTOL)]
[Code (Online CBF Adapation)]
I INTRODUCTION
Autonomous systems must operate while respecting safety constraints such as remaining within operational bounds or avoiding collisions. Control Barrier Functions have emerged as an effective tool for enforcing such constraints by ensuring forward invariance of a predefined safe set [1]. Enforcing CBF constraints for safety defines a class of optimization problems, typically Quadratic Programs and Model Predictive Control (MPC) formulations, that can be solved efficiently in real time [2, 3]. This property has made CBFs popular in a range of applications, from robotic navigation [4] to multi-agent coordination [5].
However, the definition of CBF constraints entails a class of functions called class $\mathcal{K}_{\infty}$ functions, which we refer to as CBF parameters. Finding or tuning the CBF parameters is particularly challenging for highly nonlinear and complex systems and becomes even more difficult when input constraints are present [6]. In practice, one can resort to fixing the CBF parameters a priori based on trial and error. However, using manually-tuned parameter values can lead to overly conservative motions or unsafe behavior when the environment or operating conditions change.
In this paper, we propose a novel framework for online adaptation of CBF parameters to overcome these limitations. Instead of using a fixed set of parameters for all states and all time, we allow these parameters to vary dynamically with the system state, provided they are validated locally to maintain safety over a finite time horizon. We begin by formally defining the concept of a locally validated CBF parameter and establishing theoretical conditions under which adapting these parameters ensures that the system trajectories remain safe, i.e., within the safety set over a finite horizon. The theoretical foundations of our framework are developed using the concept of the tangent cone to a set, utilized in Nagumo’s theorem for set invariance, which enable the controller to enforce safety constraints that are tailored to the current state and local operating conditions, rather than relying on uniformly worst-case assumptions.
Figure 1: Conceptual illustration of the inner safe set and locally validated CBF parameter. (a) Candidate inner safe set $\mathcal{C}^{*}(\tilde{\bm{\alpha}})$ defined via an ICCBF with a CBF parameter $\tilde{\bm{\alpha}}$. At boundary point ${\bm{x}}_{1}$, the CBF-constrained dynamics $G({\bm{x}}_{1};\tilde{\bm{\alpha}})$ is non-empty, whereas at ${\bm{x}}_{2}$ it is empty, indicating that the set cannot be rendered forward invariant. (b) With locally validated CBF parameters, the trajectory remains within the inner safe set shown in (a) over the finite-horizon, ensuring safety for that interval. By adapting the CBF parameters, the corresponding inner safe set is reshaped dynamically, alleviating conservatism by allowing the trajectory to extend beyond a fixed, globally verified inner safe set $\mathcal{C}^{*}(\bm{\alpha})$.
Several recent works have recognized the need to adjust CBF parameters online. For instance, in [7, 8], these parameters are optimized to maintain controller feasibility at each time step; however, these approaches fall short of ensuring that safety is maintained for all time. Other methodologies leverage neural networks or reinforcement learning policies to dynamically adapt CBF parameters. Although these techniques have demonstrated promising performance in complex environments, their safety guarantees often rely on unrealistic assumptions, such as unbounded control authority [9] or perfect model predictions [10, 11, 12]. When those assumptions are relaxed, safety can no longer be guaranteed.
A critical factor in the reliability of learning-based adaptation and control is the treatment of uncertainty in neural network predictions. Neural network uncertainty can be broadly classified into two categories: aleatoric and epistemic [13]. Aleatoric uncertainty stems from inherent data stochasticity such as observation and process noise, whereas epistemic uncertainty arises from insufficient training data. Neglecting these uncertainties can lead to unreliable predictions, particularly when facing out-of-distribution inputs, which may lead to catastrophic failures in safety-critical applications.
Motivated by these findings, we propose a learning-based approach grounded in the previously described theoretical framework for online adaptation of CBF parameters. In particular, we extend our previous work, which employed Probabilistic Ensemble Neural Networks (PENNs) [14], and integrate it into our framework as one concrete solution for adapting CBF parameters. Our approach explicitly accounts for both epistemic and aleatoric uncertainties inherent in neural network predictions, ensuring that the adapted parameters consistently satisfy the theoretical conditions required for local safety guarantees.
To demonstrate the scalability and effectiveness of the proposed framework, we apply it to a Vertical Takeoff and Landing (VTOL) quadplane model during transition and landing maneuvers. These scenarios are particularly challenging for traditional CBF methods due to the high dimensionality and inherent nonlinearities of the aircraft dynamics. Our simulation results show that the proposed method successfully maintains safety and enhances performance compared to fixed-parameter approaches.
In summary, this paper makes the following contributions:
  • We propose a novel framework for online CBF parameter adaptation by introducing the concept of locally validated CBF parameters, whose theoretical foundations of safety are based on tangent cone (Sec. III).
  • We raise common pitfalls in existing heuristic and learning-based CBF parameter adaptation methods from the recent literature and explain why they may fail to guarantee safety. (Sec. IV).
  • We propose a learning-based approach to adapt CBF parameters online with verification that explicitly accounts for neural network uncertainty, which adheres to our theoretical framework (Sec. V).
  • We apply our method to a VTOL quadplane transition and landing scenario. To the best of our knowledge, this is the first application of CBFs to VTOL quadplane control tasks, demonstrating its applicability to highly nonlinear systems with reduced conservatism (Sec. VI).
II BACKGROUND INFORMATION
II-A Tangent Cone
Consider a continuous-time nonlinear system:
$\dot{{\bm{x}}}=f({\bm{x}})+g({\bm{x}}){\bm{u}},$ (1)
where ${\bm{x}}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the state and ${\bm{u}}\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input, with $\mathcal{U}$ being the set of admissible controls for System (1). We assume that the functions $f:\mathcal{X}\to\mathbb{R}^{n}$ and $g:\mathcal{X}\to\mathbb{R}^{n\times m}$ are locally Lipschitz continuous. Under a chosen control law ${\bm{u}}=\pi({\bm{x}})$, which is assumed to assure that the solution to the closed-loop system exists and is unique, the closed-loop dynamics can be written as:
$\dot{{\bm{x}}}=f_{\textup{cl}}({\bm{x}})\coloneqq f({\bm{x}})+g({\bm{x}})\pi({\bm{x}}).$ (2)
Definition 1 (Tangent Cone [15]).
Let $\mathcal{S}\subset\mathbb{R}^{n}$ be a closed set. The tangent cone to $\mathcal{S}$ at ${\bm{x}}\in\mathcal{S}$ is defined as
$\mathcal{T}_{\mathcal{S}}({\bm{x}})\coloneqq\left\{{\bm{z}}\in\mathbb{R}^{n}\mid\liminf_{\tau\to 0}\frac{\textup{dist}({\bm{x}}+\tau{\bm{z}},\mathcal{S})}{\tau}=0\right\}.$ (3)
[…]
II-B Safety Analysis of CBFs via Tangent Cone
First, we demonstrate how CBFs guarantee safety through the notion of the tangent cone. Nagumo’s Theorem provides a necessary and sufficient condition for the forward invariance of a set.
Definition 2 (Forward Invariance).
A set $\mathcal{S}\subset\mathbb{R}^{n}$ is rendered forward invariant by a feedback controller $\pi:\mathcal{S}\to\mathcal{U}$ for the closed-loop system (2) if for all ${\bm{x}}(0)\in\mathcal{S}$, the solution ${\bm{x}}(t)\in\mathcal{S}$ for all $t\geq t_{0}$.
Theorem 1 (Nagumo’s Theorem [16]).
Consider the closed-loop system (2) [^2] While original Nagumo’s theorem was stated for unforced systems, here it has been extended to the closed-loop controlled system. and let $\mathcal{S}\subset\mathbb{R}^{n}$ be a closed set. Then, $\mathcal{S}$ is rendered forward invariant by $\pi$ if and only if
$f({\bm{x}})+g({\bm{x}})\pi({\bm{x}})\in\mathcal{T}_{\mathcal{S}}({\bm{x}})\quad\forall{\bm{x}}\in\mathcal{S}.$ (4)
Since $\mathcal{T}_{\mathcal{S}}({\bm{x}})=\mathbb{R}^{n}$ for any ${\bm{x}}\in\textup{Int}(\mathcal{S})$, it is sufficient to verify the condition on $\partial\mathcal{S}$.
The safety guarantee provided by a CBF can be interpreted through the lens of the tangent cone as characterized by Nagumo’s Theorem. For extra clarity in this paper, we first introduce the concept of a candidate CBF. From here, we denote the set $\mathcal{S}$ as the set of instantaneously safe states, which can often be encoded by a set of hard constraints provided to the system as requirements.
Definition 3 (Candidate CBF).
Let $\tilde{h}:\mathcal{X}\to\mathbb{R}$ be a continuously differentiable function. Any state-based constraint function $\tilde{h}$ that defines the set $\mathcal{C}\coloneqq\{{\bm{x}}\in\mathcal{X}\mid\tilde{h}({\bm{x}})\geq 0\}$, where $\mathcal{C}\subseteq\mathcal{S}$, is called a candidate CBF.
A candidate CBF becomes a CBF when it is paired with a properly chosen extended class $\mathcal{K}_{\infty}$ function $\alpha$. More precisely, we have the following definition:
Definition 4 (CBF [1]).
Let $h:\mathcal{X}\to\mathbb{R}$ be a continuously differentiable function and define the set $\mathcal{C}\coloneqq\{{\bm{x}}\in\mathcal{X}\mid h({\bm{x}})\geq 0\}$. The function $h$ is a CBF on $\mathcal{C}$ for System (2) if there exists an extended class $\mathcal{K}_{\infty}$ function $\alpha$ such that
$\sup_{{\bm{u}}\in\mathcal{U}}\left[L_{f}h({\bm{x}})+L_{g}h({\bm{x}}){\bm{u}}\right]\geq-\alpha(h({\bm{x}}))\quad\forall{\bm{x}}\in\mathcal{C},$ (5)
where $L_{f}h({\bm{x}})$ and $L_{g}h({\bm{x}})$ denote the Lie derivatives of $h$ along $f$ and $g$, respectively.
Given a CBF $h$ and its corresponding function $\alpha$, the set of all control inputs that render $\mathcal{C}\subseteq\mathcal{S}$ forward invariant, thus ensuring the system remains safe, is defined by the CBF constraint:
$K_{\textup{cbf}}({\bm{x}};\alpha)\coloneqq\left\{{\bm{u}}\in\mathcal{U}\mid L_{f}h({\bm{x}})+L_{g}h({\bm{x}}){\bm{u}}\geq-\alpha(h({\bm{x}}))\right\}.$ (6)
We introduce two set-valued maps as follows. For each state ${\bm{x}}\in\mathcal{X}$, define
$F({\bm{x}})\coloneqq\left\{f({\bm{x}})+g({\bm{x}}){\bm{u}}\mid{\bm{u}}\in\mathcal{U}\right\}$ (7)
which describes all feasible state derivatives under any admissible control. Define also
$G({\bm{x}};\alpha)\coloneqq\left\{f({\bm{x}})+g({\bm{x}}){\bm{u}}\mid{\bm{u}}\in K_{\textup{cbf}}({\bm{x}};\alpha)\right\}$ (8)
which captures the state derivatives that result from control inputs that satisfy the CBF constraint. Clearly, $G({\bm{x}};\alpha)$ is a safety-restricted subset of $F({\bm{x}})$.
The following lemma establishes the connection between the CBF condition and forward invariance of the set $\mathcal{C}$ via the tangent cone.
Lemma 1.
Consider the system:
$\dot{{\bm{x}}}(t)\in G({\bm{x}};\alpha).$ (9)
Suppose that $h$ is a CBF for System (2) with a given $\alpha$, and let $\mathcal{C}$ be defined as in Definition 4. If ${\bm{x}}(\cdot)$ is any trajectory of (9) under a control policy satisfying ${\bm{u}}(t)\in K_{\textup{cbf}}\big{(}{\bm{x}}(t);\alpha\big{)}$ with ${\bm{x}}(0)\in\mathcal{C}$, then ${\bm{x}}(t)\in\mathcal{C}$ for all $t\geq t_{0}$.
Proof.
Since $h$ is a CBF, the CBF condition (5) guarantees
$G({\bm{x}};\alpha)\subseteq\mathcal{T}_{\mathcal{C}}({\bm{x}})\quad\forall{\bm{x}}\in\partial\mathcal{C}.$ (10)
By Theorem 1, the set $\mathcal{C}$ is forward invariant. ∎
II-C Input Constrained CBFs
We extend the notion of CBFs to incorporate input constraints, yielding Input Constrained CBFs (ICCBFs) [6], a generalization of High Order CBFs (HOCBFs) [17].
Assume that the original function $h$ is an $r^{\textup{th}}$ order differentiable function. We define a series of functions $b_{i}:\mathcal{X}\to\mathbb{R}$, $i=0,\ldots,r-1$, as
$\begin{split}b_{0}({\bm{x}};\bm{\alpha})&\coloneqq h({\bm{x}}),\\ b_{1}({\bm{x}};\bm{\alpha})&\coloneqq\inf_{{\bm{u}}\in\mathcal{U}}\left[\dot{b}_{0}({\bm{x}},{\bm{u}};\bm{\alpha})\right]+\alpha_{1}(b_{0}({\bm{x}};\bm{\alpha})),\\ &\vdots\\ b_{r-1}({\bm{x}};\bm{\alpha})&\coloneqq\inf_{{\bm{u}}\in\mathcal{U}}\left[\dot{b}_{r-2}({\bm{x}},{\bm{u}};\bm{\alpha})\right]+\alpha_{r-1}(b_{r-2}({\bm{x}};\bm{\alpha})),\end{split}$ (11)
and the constraint function $b_{r}:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ as
$b_{r}({\bm{x}},{\bm{u}};\bm{\alpha})\coloneqq\dot{b}_{r-1}({\bm{x}},{\bm{u}};\bm{\alpha})+\alpha_{r}\left(b_{r-1}({\bm{x}};\bm{\alpha})\right),$ (12)
where $\bm{\alpha}=\{\alpha_{1},\ldots,\alpha_{r}\}$ is a set of class $\mathcal{K}$ functions. We also define the following sets:
$\begin{split}\mathcal{C}_{0}(\bm{\alpha})&\coloneqq\{{\bm{x}}\in\mathcal{X}\mid b_{0}({\bm{x}};\bm{\alpha})\geq 0\}\subseteq\mathcal{S}\\ \mathcal{C}_{1}(\bm{\alpha})&\coloneqq\{{\bm{x}}\in\mathcal{X}\mid b_{1}({\bm{x}};\bm{\alpha})\geq 0\}\\ &\vdots\\ \mathcal{C}_{r-1}(\bm{\alpha})&\coloneqq\{{\bm{x}}\in\mathcal{X}\mid b_{r-1}({\bm{x}};\bm{\alpha})\geq 0\}.\end{split}$ (13)
Definition 5 (Inner Safe Set [14]).
The set $\mathcal{C}^{*}(\bm{\alpha})\subseteq\mathcal{S}$,
$\mathcal{C}^{*}(\bm{\alpha})\coloneqq\bigcap_{i=0}^{r-1}\mathcal{C}_{i}(\bm{\alpha}),$ (14)
where $\mathcal{C}_{i}(\bm{\alpha})$ is defined in (13) is called an inner safe set for System (2).
Definition 6 (ICCBF [6]).
The function $h$ is a ICCBF [^3] In [6], $b_{r-1}$ is referred to as ICCBF; in this paper, for notational consistency with HOCBF [17], we designate $h$ as ICCBF, with all corresponding theoretical properties preserved. on $\mathcal{C}^{*}(\bm{\alpha})$ for System (2) if there exist class $\mathcal{K}$ functions $\bm{\alpha}=\{\alpha_{1},\ldots,\alpha_{r}\}$ such that
$\sup_{{\bm{u}}\in\mathcal{U}}\left[b_{r}({\bm{x}},{\bm{u}};\bm{\alpha})\right]\geq 0\quad\forall{\bm{x}}\in\mathcal{C}^{*}(\bm{\alpha}).$ (15)
This condition allows the formulation of safety constraints for systems of relative degree at most $r$.
Analogously to CBF in Sec. II-B, we establish the connection between ICCBFs and Nagumo’s theorem via the tangent cone. The set of admissible control inputs that renders the inner safe set $\mathcal{C}^{*}(\bm{\alpha})\subseteq\mathcal{S}$ forward invariant is defined as
$K_{\textup{iccbf}}({\bm{x}};\bm{\alpha})\coloneqq\left\{{\bm{u}}\in\mathcal{U}\mid b_{r}({\bm{x}},{\bm{u}};\bm{\alpha})\geq 0\right\}.$ (16)
The set of allowable state derivatives is given by
$G({\bm{x}};\bm{\alpha})\coloneqq\left\{f({\bm{x}})+g({\bm{x}}){\bm{u}}\mid{\bm{u}}\in K_{\textup{iccbf}}({\bm{x}};\bm{\alpha})\right\}.$ (17)
Lemma 2.
Given an ICCBF $h$ with a specified $\bm{\alpha}$, suppose the system evolves according to
$\dot{{\bm{x}}}(t)\in G({\bm{x}};\bm{\alpha}).$ (18)
Then, the ${\bm{x}}(t)\in\mathcal{C}^{*}(\bm{\alpha})$ for all $t\geq t_{0}$.
Proof.
By the same principle as in Lemma 1, we have
$G({\bm{x}};\bm{\alpha})\subseteq\mathcal{T}_{\mathcal{C}^{*}(\bm{\alpha})}({\bm{x}})\quad\forall{\bm{x}}\in\partial\mathcal{C}^{*}(\bm{\alpha}),$ (19)
which implies that $\mathcal{C}^{*}(\bm{\alpha})$ is forward invariant. ∎
III METHOD OVERVIEW
A widely recognized limitation of CBFs is the difficulty in finding the extended class $\mathcal{K}_{\infty}$ function in (5). This challenge becomes even more pronounced in the formulation of ICCBFs, because (a) multiple class $\mathcal{K}_{\infty}$ functions $\bm{\alpha}=\{\alpha_{1},\ldots,\alpha_{r}\}$ must be tuned simultaneously in (15), and (b) they also determine the inner safe set $\mathcal{C}^{*}(\bm{\alpha})$. In what follows, the extended class $\mathcal{K}_{\infty}$ functions $\bm{\alpha}$ will be referred to as the CBF parameter.
Remark 1.
Given a candidate ICCBF [^4] Similar to candidate CBFs in Definition 3, candidate ICCBFs are constructed following the ICCBF approach, but the ICCBF condition has not yet been verified with appropriate CBF parameters. $\tilde{h}$ with a specified parameter $\tilde{\bm{\alpha}}$, we define the set of control inputs satisfying the candidate ICCBF constraint as
$K_{\textup{cand}}({\bm{x}};\tilde{\bm{\alpha}})\coloneqq\left\{{\bm{u}}\in\mathcal{U}\mid b_{r}({\bm{x}},{\bm{u}};\tilde{\bm{\alpha}})\geq 0\right\}.$ (20)
There exist values of $\tilde{\bm{\alpha}}$ for which, at some states ${\bm{x}}\in\partial\mathcal{C}^{*}(\tilde{\bm{\alpha}})$ (see Fig. 1a), the set $K_{\textup{cand}}({\bm{x}};\tilde{\bm{\alpha}})$ is empty, resulting the corresponding set $G({\bm{x}};\tilde{\bm{\alpha}})$ is also empty. This implies that the inner safe set $\mathcal{C}^{*}(\tilde{\bm{\alpha}})$ cannot be rendered forward invariant by any control law satisfying the candidate ICCBF constraint.
To this end, this paper presents the following problem:
Problem 1 (CBF Parameter Adaptation Framework).
Given the closed-loop system (2) and a safe set $\mathcal{S}$ (for which a valid ICCBF is unknown), design a CBF parameter adaptation framework that yields adapted parameter $\tilde{\bm{\alpha}}(t)$ such that a CBF-based controller satisfying
${\bm{u}}(t)\in K_{\textup{cand}}({\bm{x}}(t);\tilde{\bm{\alpha}}(t)),$ (21)
where $K_{\textup{cand}}$ is given by (20), guarantees that ${\bm{x}}(t)\in\mathcal{S}$ for all $t\geq t_{0}$.
To address Problem 1, we propose a new concept of locally validated CBF parameters (see Fig. 1b), which are designed to guarantee safety over a finite time interval, i.e., ${\bm{x}}(\tau)\in\mathcal{S}$ for all $\tau\in[0,T]$, where $T<\infty$. We denote ${\bm{x}}_{t+\tau|t}$ as the predicted state at time $t+\tau$ computed at time $t$, and ${\bm{u}}_{t+\tau|t}=\pi({\bm{x}}_{t+\tau|t})$ for the corresponding control input along the trajectory.
Definition 7 (Locally Validated CBF Parameter [^5] It was also termed a locally valid parameter in [14].).
Let $t\geq t_{0},\tau\geq 0$ and $T>0$. Denote by ${\bm{x}}_{t+\tau|t}$ the predicted state with initial condition ${\bm{x}}_{t|t}\in\mathcal{C}^{*}(\tilde{\bm{\alpha}}(t))$, and let $\pi({\bm{x}}_{t+\tau|t})\in K_{\textup{cand}}({\bm{x}}_{t+\tau|t};\tilde{\bm{\alpha}}(t))$. The CBF parameter $\tilde{\bm{\alpha}}(t)$ is said to be locally validated over the time interval $[t,t+T]$ if, for all $\tau\in[0,T]$,
$f({\bm{x}}_{t+\tau|t})+g({\bm{x}}_{t+\tau|t})\pi({\bm{x}}_{t+\tau|t})\in\mathcal{T}_{\mathcal{C}^{*}(\tilde{\bm{\alpha}}(t))}({\bm{x}}_{t+\tau|t}).$ (22)
Similar to Theorem 1, this condition is only non-trivial on $\partial\mathcal{C}^{*}(\tilde{\bm{\alpha}}(t))$.
Theorem 2 (Main Result).
Consider the closed-loop system (2) with control inputs ${\bm{u}}\in K_{\textup{cand}}({\bm{x}};\tilde{\bm{\alpha}}(t))$. Suppose that the system is initialized with a locally validated CBF parameter $\tilde{\bm{\alpha}}(t_{0})$ at $t=t_{0}$, and that it is adapted at times $(t_{k})_{k\in\mathbb{N}\cup\{0\}}$ with $t_{k+1}-t_{k}<T$ for all $k$ such that $\tilde{\bm{\alpha}}(t_{k})$ satisfies (22) over $[t_{k},t_{k}+T]$. Then, the system trajectories remain within $\mathcal{S}$ for all $t\geq t_{0}$.
Proof.
We prove this by showing that for all $k\in\mathbb{N}\cup\{0\}$, the state remains in $\mathcal{S}$ over the interval $[t_{k},t_{k}+T]$. The proof is by induction.
Base Case: At $t_{k=0}=t_{0}$, the system is initialized with a locally validated CBF parameter $\tilde{\bm{\alpha}}(t_{0})$. By Definition 7, the closed-loop system (2) with control inputs chosen from $K_{\textup{cand}}({\bm{x}};\tilde{\bm{\alpha}}(t_{0}))$ satisfies the conditions of Nagumo’s Theorem over the finite interval $[0,T]$. Consequently, ${\bm{x}}(t)\in\mathcal{C}^{*}(\tilde{\bm{\alpha}}(t_{0}))\subseteq\mathcal{S}$, $\forall t\in[t_{0},t_{0}+T]$.
Inductive Step: Suppose the claim is true for some $k>0$, i.e., the state remains in $\mathcal{C}^{*}(\tilde{\bm{\alpha}}(t_{k}))\subseteq\mathcal{S}$ for all $t\in[t_{k},t_{k}+T]$. By the adaptation, the CBF parameter is updated at some time $t_{k+1}$ such that it is locally validated over $[t_{k+1},t_{k+1}+T]$, where $t_{k+1}<t_{k}+T$. Again, by Definition 7 and Nagumo’s Theorem, the system is ensured to remain in $\mathcal{C}^{*}(\tilde{\bm{\alpha}}(t_{k+1}))\subseteq\mathcal{S}$ over $[t_{k+1},t_{k+1}+T]$. ∎
A significant advantage of the proposed CBF parameter adaptation framework is twofold. First, it allows candidate ICCBFs $\tilde{h}$ to be used without tuning the fixed CBF parameter $\bm{\alpha}$ for deployment. Second, it alleviates the inherent conservatism of globally validated parameter $\bm{\alpha}$. A conservative CBF parameter typically confines the system to an inner safe set $\mathcal{C}^{*}(\bm{\alpha})$ that is overly restrictive. In contrast, by verifying and adapting locally validated CBF parameter $\tilde{\bm{\alpha}}(t)$, the framework effectively reshapes the inner safe set $\mathcal{C}^{*}(\tilde{\bm{\alpha}}(t))$ in response to the current state ${\bm{x}}_{t|t}$ and nearby operating conditions. This permits the system trajectory to extend beyond the conservative inner safe set while ensuring safety over the finite interval. For the remainder of the paper, we drop the argument $t$ whenever clear from the context.
Now, we narrow our focus to the key sub-problem:
Problem 2 (Identifying Locally Validated CBF Parameters).
[…]
The subsequent sections (Sec. IV-Sec. V) address possible solutions to Problem 2.
IV PRIOR LEARNING-BASED APPROACHES TO CBF PARAMETER ADAPTATION
The core idea behind solving Problem 1 is to identify online locally validated CBF parameters $\tilde{\bm{\alpha}}(t_{k})$, $k\in\mathbb{N}\cup\{0\}$. A naïve approach is to resort to an exhaustive search by forward propagating the closed-loop system trajectory over all possible parameters; however, this method is computationally intractable. An alternative is to collect data samples offline and leverage prediction models for online estimation. Neural networks, as universal function approximators, have demonstrated exceptional capability in handling large training datasets while providing constant-time inference, making them attractive candidates for this task.
A spontaneous solution might be to train a neural network to predict the optimal CBF parameters online and then directly adapt those parameters. Let us denote the input by $\bar{X}=[{\bm{x}}_{t|t}^{\top},\mathcal{E}_{t}^{\top}]^{\top}$, where ${\bm{x}}_{t|t}$ is the system state and $\mathcal{E}_{t}$ represents additional environmental information (e.g., obstacle information in a collision avoidance task) at time $t$, and denote the output by $\bar{Y}=\tilde{\bm{\alpha}}^{*}$ representing the optimal parameter $\tilde{\bm{\alpha}}^{*}\in\mathcal{A}$. Then, a deterministic neural network ${\bf E}$, parameterized by $\bm{\theta}$, defines the mapping (see Fig. 2a):
$\tilde{\bm{\alpha}}^{*}={\bf E}(\bar{X};\bm{\theta}).$ (23)
While the works in [9, 10, 11, 12] do not explicitly implement our locally validated CBF parameter adaptation framework, they fall into this category and have demonstrated promising results in reducing conservatism through online adaptation of CBF parameters using neural networks. However, these approaches do not sufficiently account for the inherent uncertainty in neural network predictions. Consequently, the safety of the system cannot be guaranteed, as the predicted CBF parameter $\tilde{\bm{\alpha}}^{*}$ out of (23) might not lie within the set of locally validated CBF parameters $\mathcal{A}$ given by Problem 2.
V UNCERTAINTY-AWARE ADAPTATION OF LOCALLY VALIDATED CBF PARAMETERS
The approach in the previous section, as illustrated by (23), essentially treats the trained model as infallible, offering no mechanism to judge the confidence of those predictions. In contrast, we propose a probabilistic verification method for Problem 2 as shown in Fig. 2: instead of outputting a single set of CBF parameters $\tilde{\bm{\alpha}}^{*}$, our proposed method takes a candidate set of CBF parameters as input and predicts the distributions of a safety margin $\phi$ and a performance metric $\delta$, which are defined later on. The key novelty is that this approach enables us to rigorously verify locally validated CBF parameters despite uncertainty in neural network predictions. This way, the burden of ensuring safety is shifted from the trained model to a verification process that interprets its output with uncertainty quantification.
V-A Probabilistic Ensemble Neural Network Model
We employ a Probabilistic Ensemble Neural Network (PENN) as the prediction model, denoted as ${\bf F}$, to capture both prediction uncertainty and inherent output variability [18]. Building on the input $\bar{X}$ from Sec. IV, we augment it with the CBF parameter of interest $\tilde{\bm{\alpha}}$. Given this augmented input $X=[\bar{X}^{\top},\tilde{\bm{\alpha}}]^{\top}$, the model outputs two quantities $Y=[\phi,\delta]^{\top}$: a safety margin $\phi\in\mathbb{R}$ and a performance metric $\delta\in\mathbb{R}$ (see Fig. 2b). Concretely, we define $\phi$ as the minimum of the ICCBF constraint function over the interval $[t,t+T]$:
$\phi\coloneqq\min_{\tau\in[0,T]}\,b_{r}\left({\bm{x}}_{t+\tau|t},\,{\bm{u}}_{t+\tau|t};\,\tilde{\bm{\alpha}}\right).$ (24)
Notably, this minimum value is sufficient to verify the local validity condition because, as stated in Definition 7, the condition (22) is only non-trivial on the boundary of the inner safe set $\partial\mathcal{C}^{*}(\tilde{\bm{\alpha}})$. If $\phi$ remains above zero, then $K_{\textup{cand}}({\bm{x}};\tilde{\bm{\alpha}})$ is non-empty throughout $[t,t+T]$, ensuring that the system stays within $\mathcal{C}^{*}(\tilde{\bm{\alpha}})$ over that interval. Meanwhile, $\delta$ reflects the controller’s performance under the CBF parameter $\tilde{\bm{\alpha}}$ (e.g., capturing the rate at which the robot progresses toward a goal); we will discuss its criteria and usage in Sec. V-C.
Offline Training: We train an ensemble of $B$ neural networks, $\{{\bf E}_{1},\dots,{\bf E}_{B}\}$, $B\geq 2$, where each network parameterized by $\bm{\theta}_{b}$ estimates the mean ${\bm{\mu}}_{\bm{\theta}_{b}}(X)$ and diagonal covariance ${\bm{\Sigma}}_{\bm{\theta}_{b}}(X)$ of a Gaussian distribution. Each member is trained using a Gaussian negative log-likelihood loss. The predictions of all ensemble members are then combined to form a Gaussian Mixture Model (GMM) that represent the PENN’s belief over the output:
$\hat{Y}\sim{\bf F}(X;\bm{\theta}_{1:B})=\frac{1}{B}\sum_{b=1}^{B}{\,\mathcal{N}\left({\bm{\mu}}_{\bm{\theta}_{b}}(X),{\bm{\Sigma}}_{\bm{\theta}_{b}}(X)\right)},$ (25)
as we use equal weights for each member.
To capture epistemic uncertainty, we independently initialize each ensemble member with random weights $\bm{\theta}_{b}$. The resulting disagreement among ensemble predictions serves as a quantitative measure of epistemic uncertainty. Another notable advantage is that the forward pass of the PENN for prediction can be fully parallelized across ensemble members, which is beneficial for real-time implementation [19].
Figure 2: Neural network configurations for CBF parameter adaptation. (a) Prior approaches: A deterministic neural network directly outputs the optimal CBF parameter $\tilde{\bm{\alpha}}^{*}$. (b) Proposed method: A PENN model augments the input with the CBF parameter $\tilde{\bm{\alpha}}$ of interest, and outputs a GMM distribution representing the predicted characteristics when that CBF parameter is chosen, thereby enabling uncertainty-aware verification.
V-B Online Uncertainty-Aware Verification
We now present a verification process to assess whether the given CBF parameter is locally validated.
V-B1 Verification Under Epistemic Uncertainty
In this step, we first evaluate the epistemic uncertainty associated with the PENN predictions. Specifically, we quantify the disagreement among ensemble members using the Jensen-Rényi divergence (JRD) with quadratic Rényi entropy, which admits a closed-form expression for Gaussian mixtures [20]. In essence, the JRD $D(X;{\bf E}_{1:B})\in\mathbb{R}^{\dim(Y)}$ measures how “spread out” or inconsistent the ensemble predictions are.
If the computed JRD exceeds a predefined threshold $D_{\textup{thr}}\in\mathbb{R}^{\dim(Y)}$, the corresponding input is deemed to be out-of-distribution and its prediction is discarded. This serves as our first “gate” to ensure that only high-confidence predictions interpolating within the training data, rather than extrapolating far beyond them, are used to determine the local validity of the CBF parameter.
V-B2 Verification Under Aleatoric Uncertainty
For prediction results that pass the epistemic filter, we next examine the predicted distribution of the safety margin $\phi$ to ensure that the local validity condition is maintained with high probability despite inherent aleatoric uncertainties. Recall that $\phi\geq 0$ in (24) implies that the given CBF parameter $\tilde{\bm{\alpha}}$ is locally validated. For notational simplicity, we treat the safety margin $\phi$ as the only output in $Y$ throughout this subsection, and we denote the predicted safety margin as $\hat{\phi}$.
We utilize the Conditional Value at Risk (CVaR) to assess the left-tail risk associated with $\hat{\phi}$ [21]. For a single ensemble member’s prediction $\hat{\phi}\sim{\bf E}_{b}(X)$, the left Value at Risk (VaR) at confidence level $\varepsilon$ is defined as
$\textnormal{VaR}_{\varepsilon}^{{\bf E}_{b}(X)}\left(\hat{\phi}\right)\coloneqq\sup_{\nu\in\mathbb{R}}\left\{\nu\,|\,\textnormal{Prob}^{{\bf E}_{b}(X)}\left(\hat{\phi}<\nu\right)\leq\varepsilon\right\}.$ (26)
Intuitively, $\textup{VaR}_{\varepsilon}$ represents the worst-case safety margin that is exceeded with probability at least $1-\varepsilon$.
Building on VaR, we define CVaR using an optimization formulation as proposed in [22]:
$\textnormal{CVaR}_{\varepsilon}^{{\bf E}_{b}(X)}\left(\hat{\phi}\right)\coloneqq\sup_{\eta\in\mathbb{R}}\left\{\eta-\frac{1}{\varepsilon}\mathbb{E}_{{\bf E}_{b}(X)}\left[(\eta-\hat{\phi})^{+}\right]\right\},$ (27)
where $(\cdot)^{+}\coloneqq\max\{\cdot,0\}$. The left CVaR is interpreted as the expected safety margin that falls in the left tail of the distribution, covering the worst $\varepsilon$-fraction of cases.
To robustly account for aleatoric uncertainty, we treat the ensemble set $\mathfrak{E}=\{{\bf E}_{1},\ldots,{\bf E}_{B}\}$ as an ambiguity set and impose the following distributionally robust CVaR constraint [23]:
$\sup_{{\bf E}_{b}\in\mathfrak{E}}\textnormal{CVaR}{\varepsilon}^{{\bf E}_{b}(X)}\left(\hat{\phi}\right)\geq 0.$ (28)
This condition is equivalent to requiring that
$\inf_{{\bf E}_{b}\in\mathfrak{E}}\textnormal{Prob}^{{\bf E}_{b}(X)}\left(\hat{\phi}\geq 0\right)\geq 1-\varepsilon,$ (29)
i.e., the given $\tilde{\bm{\alpha}}$ is accepted only if the predicted safety margin $\hat{\phi}$ is non-negative with probability at least $1-\varepsilon$.
In summary, after filtering out predictions with high epistemic uncertainty, the distributionally robust CVaR constraint ensures probabilistically that the local validity condition (22) is met under aleatoric uncertainties. This two-step verification process yields a verified set of locally validated CBF parameters:
[…]
where $\mathcal{A}_{\textup{cand}}$ denotes the batch of candidate CBF parameters.
In implementation, we perform batch-wise inference on $\mathcal{A}_{\textup{cand}}$ with the PENN model, benefiting from GPU parallelization. Moreover, since both JRD and CVaR admit analytic solutions under the normal distribution assumption, the uncertainty quantification can also be computed entirely in parallel on GPU. We emphasize that there is no single iteration loop in the entire CBF parameter verification and adaptation procedure, ensuring efficient real-time implementation.
V-C Online CBF Parameter Adaptation
From the verified set $\mathcal{A}$ (30), we select the final CBF parameter based on a performance metric $\delta\in Y$. Here, $\delta$ is defined as the accumulated time over the forward propagation interval $[t,t+T]$ during which the system’s speed towards the goal falls below a specified threshold, with lower values indicating more efficient progress. Hence, we adapt the CBF parameters by choosing: $\tilde{\bm{\alpha}}^{*}=\operatorname*{arg\,min}_{\tilde{\bm{\alpha}}\in\mathcal{A}}\hat{\delta}$, where $\hat{\delta}$ is the predicted performance metric associated with the candidate parameter $\tilde{\bm{\alpha}}$. This yields a less restrictive CBF-constrained control input set $K_{\textup{cand}}({\bm{x}};\tilde{\bm{\alpha}}^{*})$, thereby improving performance.
VI RESULTS
Figure 3: Illustration of a VTOL quadplane aircraft.
Figure 4: We visualize the aircraft trajectory along with its speed profile, rigid-body pose, and elevator angle $\delta_{e}$, with obstacles depicted in gray. (a) With fixed low parameters, the trajectory shows a significant altitude detour as the ICCBF constraints force the aircraft to pitch up to decelerate in response to the obstacle. (b) With fixed high parameters, the controller becomes infeasible, ultimately resulting in a collision. (c) Our adaptive approach dynamically adjusts the CBF parameters based on the aircraft’s speed and position. Initially, due to the high speed, it maintains low CBF parameters, prompting the elevator to pitch up and generate additional drag. As the aircraft slows down, the parameters increase to enhance performance, enabling the aircraft to fly safely beneath the obstacle and executing a smooth transition afterwards. Inset: Evolution of the adapted CBF parameters over time.
VI-A VTOL Quadplane
We consider the 2D VTOL quadplane, illustrated in Fig. 3, as our simulation case study. This platform exhibits complex, high-dimensional nonlinear dynamics and strict input constraints, posing significant challenges for safety-critical control design, particularly during transitions between flight modes. Tuning the parameters of CBFs is a very challenging task for this class of dynamics.
The state is defined as ${\bm{x}}=[{\bm{p}}^{\top},{\bm{v}}^{\top},\theta,q]^{\top}$, where ${\bm{p}}=[x,z]^{\top}\in\mathbb{R}^{2}$ denotes the planar position (horizontal and vertical) in the inertial frame (with $z$ pointing upward), ${\bm{v}}\in\mathbb{R}^{2}$ is the corresponding velocity in the inertial frame, $\theta\in[-\pi,\pi]$ (in radians) is the pitch angle, and $q=\dot{\theta}$ is the pitch rate. The control input is defined as ${\bm{u}}=[\delta_{f},\delta_{r},\delta_{p},\delta_{e}]^{\top}\in\mathbb{R}^{4}$, where $\delta_{f},\delta_{r}\in[0,1]$ are the throttles for the forward and rear rotors, $\delta_{p}\in[0,1]$ is the throttle for the forward thruster (pusher), and $\delta_{e}\in[-0.5,0.5]$ (in radians) is the elevator deflection. $\ell_{f}$ and $\ell_{r}$ are the front and rear lever arms.
The rigid-body equations of motion are given by
$\begin{split}\dot{{\bm{p}}}&={\bm{v}}\\ m\dot{{\bm{v}}}&=-mg{\bm{e}}_{z}+{\bf R}(\theta){\bf F}_{\textup{total}}\\ \dot{\theta}&=q\\ J_{y}\dot{q}&=M_{\textup{total}},\end{split}$ (31)
where $m\in\mathbb{R}_{>0}$ and $J_{y}\in\mathbb{R}_{>0}$ denote the mass and pitch moment of inertia of the quadplane, respectively, $M_{\textup{total}}\in\mathbb{R}$ is the total pitching moment experienced by the aircraft, ${\bf F}\in\mathbb{R}^{2}$ is the external force in the body-fixed frame, and ${\bm{e}}_{z}=[0,1]^{\top}$. The force ${\bf F}_{\textup{total}}$ is rotated into the inertial frame via the rotation matrix ${\bf R}(\theta)\in\mathbb{SO}(2)$.
VI-A1 Aerodynamics Forces
Let $(u,w)$ be the velocity in the body frame, i.e., $[u,w]^{\top}={\bf R}^{-1}(\theta){\bm{v}}$. The airspeed is $V=\sqrt{u^{2}+w^{2}}$, and the angle of attack is given by $\alpha=\tan^{-1}(-w/u)$  [^6] Here, we abuse the notation $\alpha$ to denote the angle of attack; elsewhere, $\alpha$ denotes an extended class $\mathcal{K}_{\infty}$ function.. We denote the various aerodynamic coefficients as $C_{(\cdot)}\in\mathbb{R}$.
To capture realistic airfoil stall effects at high angles of attack, the lift coefficient $C_{L}(\alpha)$ is chosen to be a smooth nonlinear function of $\alpha$ that blends a linear lift curve at small $|\alpha|$ with a flat-plate model at large $|\alpha|$ [24]. Concretely,
$C_{L}(\alpha)=(1-\sigma(\alpha))\left(C_{L_{0}}+C_{L_{\alpha}}\alpha\right)+\sigma(\alpha)\left(2\sin\alpha\cos\alpha\right),$ (32)
where the blending (sigmoid) function is defined as
$\sigma(\alpha)=\frac{1+e^{-w_{s}(\alpha-\alpha_{s})}+e^{w_{s}(\alpha+\alpha_{s})}}{\left(1+e^{-w_{s}(\alpha-\alpha_{s})}\right)\left(1+e^{w_{s}(\alpha+\alpha_{s})}\right)},$ (33)
with $w_{s}\in\mathbb{R}_{>0}$ is the transition rate and $\alpha_{s}\in\mathbb{R}_{>0}$ is the stall angle. The drag coefficient is modeled by the quadratic function:
$C_{D}(\alpha)=C_{D_{0}}+C_{D_{\alpha^{2}}}\alpha^{2}.$ (34)
Following [25], the aerodynamic forces in the body frame are given by
${\bf F}_{0}(V,\alpha)=\frac{1}{2}\rho V^{2}S\,{\bf R}(\alpha)\begin{bmatrix}-C_{D}(\alpha)\\ C_{L}(\alpha)\end{bmatrix}.$ (35)
Here, $\rho\in\mathbb{R}_{>0}$ denotes the air density, $S\in\mathbb{R}_{>0}$ is the reference wing area, and ${\bf R}(\alpha)$ is the rotation matrix that transforms forces from the wind frame to the body frame.
VI-A2 Control Surface and Rotor Forces
The force due to the elevator $\delta_{e}$ is modeled as
${\bf F}_{\delta_{e}}(V,\alpha)\delta_{e}=\frac{1}{2}\rho V^{2}S\,{\bf R}(\alpha)\begin{bmatrix}-C_{D_{\delta_{e}}}(\alpha)\\ C_{L_{\delta_{e}}}(\alpha)\end{bmatrix}\delta_{e}$ (36)
For the rotor thrusts, we assume a linear mapping from throttle input to force. The front, rear, and forward thrusts are given by $T_{f}=k_{f}\delta_{f}$, $T_{r}=k_{r}\delta_{r}$, and $T_{p}=k_{p}\delta_{p}$, with maximum thrust coefficients $k_{(\cdot)}$.
Finally, the total force in the body frame is
${\bf F}_{\textup{total}}(V,\alpha,{\bm{u}})={\bf F}_{0}(V,\alpha)+{\bf F}_{\delta_{e}}(V,\alpha)\delta_{e}+\begin{bmatrix}T_{p}\\ T_{f}+T_{r}\end{bmatrix}.$ (37)
Similarly, the total pitching moment, $M_{\textup{total}}$, is modeled following [25].
The overall quadplane dynamics (31) are control affine and continuously differentiable with respect to both the state and control inputs.
VI-B Experimental Setup & Results
We design a MPC-CBF controller, which solves constrained optimization problem with ICCBF constraints (12). The objective function is given as a linear quadratic cost terms that drive the quadplane toward a waypoint, which is updated once the system reaches a predefined proximity. Both the discretization and control update intervals are set to $0.05$ s. The physical parameters and aerodynamic coefficient can be found in our public code repository.  [^7] https://github.com/tkkim-robot/safe_control/blob/main/robots/vtol2D.py
The test environment, illustrated in Fig. 4, features a scenario where the quadplane initiates in cruise mode with a high initial x-axis velocity, transitions to hover mode at an intermediate waypoint $(70,10)$ m, and lands at $(70,0)$ m while avoiding obstacles. The candidate ICCBF for collision avoidance is defined by the distance-based constraint: $\tilde{h}_{j}({\bm{x}})=(x-x_{\textup{obs},j})^{2}+(z-z_{\textup{obs},j})^{2}-(l_{\textup{robot}}+l_{\textup{obs},j})^{2}$, where $(x_{\textup{obs},j},z_{\textup{obs},j})$ and $l_{\textup{obs},j}$ denote the location and radius of the $j^{\textup{th}}$ obstacle, and $l_{\textup{robot}}$ is the quadplane’s safety radius. The safety margin $\phi\in Y$ is computed as the minimum across all obstacles. We use $r=2$ to construct the candidate ICCBF, consistent with the fact that $\tilde{h}_{j}$ has a relative degree of 2 for all four control inputs. As demonstrated in [6], the ICCBF thus trivially reduces to the HOCBF.
For training, we generate $22,050$ trajectory samples offline by varying the initial state ${\bm{x}}$ and the linear CBF parameters $\tilde{\bm{\alpha}}=(\tilde{\alpha}_{1},\tilde{\alpha}_{2})$ [^8] In this work, we use linear class $\mathcal{K}_{\infty}$ functions for simplicity.. The forward propagation interval $T$, representing the maximum simulation duration, is set to $15$ s to allow sufficient time for the system to land when feasible. In cases of infeasibility or collision, the safety margin $\phi$ is assigned a large negative value to clearly distinguish those from marginally feasible cases (i.e., where $0\leq\phi\ll 1$). We do not include additional environmental information in the neural network input $X$, given that obstacles are static in this test case. The PENN model ${\bf F}$ comprises $B=3$ ensemble models, each has 5 MLP layers with ReLU activation function. During training, 3% i.i.d. Gaussian noise is added to the input $X$, excluding the CBF parameters, to simulate observation and dynamics model uncertainties. Further implementation details follow [14].
For testing, we compare our approach with two baseline MPC controllers using fixed CBF parameters–one with low parameters ($\tilde{\alpha}_{1}=\tilde{\alpha}_{2}=0.05$) and one with high parameters ($\tilde{\alpha}_{1}=\tilde{\alpha}_{2}=0.35$). For our method, the CBF parameters are allowed to adapt within the range $(0.05,0.35)$, with an initial value of $0.05$, and the adaptation is performed at every MPC iteration. We set $D_{\textup{thr}}$ in (30) to the JRD value observed when inputs are completely out-of-distribution. The initial position and x-axis velocity are set to $(2,10)$ m and $20$ m/s, respectively.
The simulation results are shown in Fig. 4. In the baseline with low parameters, the controller is significantly influenced by the obstacle from the start, leading to a significant detour in altitude as the aircraft pitches upward to decelerate. In contrast, the baseline with high parameters suffers from infeasibility, ultimately resulting in a collision. Our adaptation method updates the CBF parameters based on the aircraft’s velocity and position at every time step, achieves a reduced reach time while ensuring safety. These results demonstrate the effectiveness of our proposed adaptation strategy.
VII CONCLUSION
In this paper, we proposed a novel framework for online adaptation of (input constrained) CBF parameters by introducing locally validated CBF parameters. To effectively identify these parameters online, we developed a learning-based approach leveraging a PENN model alongside an uncertainty-aware verification process addressing neural network uncertainties. Simulation results on a VTOL quadplane demonstrate enhanced performance while rigorously maintaining safety under challenging conditions.